find the exact values of all trig functions if tan theta= -4/3 and pi/2

Answer by lwsshak3(11628)

find the exact values of all trig functions if tan theta= -4/3 and pi/2
***
tan x=-4/3
you are working with a 3-4-5 right triangle in quadrant II
sin x=-4/5
cos x=3/5
csc x=-5/4
sec x=5/3
tan x=-4/3
cot x=-3/4
..
tanx x=π/2
hypotenuse of working right triangle=√(π)^2+2^2)=√(π^2+4)
sin x=π/√(π^2+4)
cos x=2/√(π^2+4)
csc x=√(π^2+4)/π
sec x=√(π^2+4)/2
tan x=π/2
cot x=2/π